OpenStudy (anonymous):

Simplify the radical:

OpenStudy (anonymous):

\[\sqrt{8}+\sqrt{50}\]

OpenStudy (anonymous):

7 sqrt2

jimthompson5910 (jim_thompson5910):

\[\sqrt{8}+\sqrt{50}\] \[\sqrt{4*2}+\sqrt{25*2}\] \[\sqrt{4}*\sqrt{2}+\sqrt{25}*\sqrt{2}\] \[2\sqrt{2}+5\sqrt{2}\] \[7\sqrt{2}\]

OpenStudy (wasiqss):

4radical 2+5radical 2=7 radical 2